OpenStudy (anonymous):

How can i get the answer 24 by using 8,8,3,3. You can use the main signs add,subtract,multiply,divide. Brain Teaser

OpenStudy (alexandervonhumboldt2):

8*3=24 ????????????/

OpenStudy (anonymous):

Do you have to use all numbers?

OpenStudy (anonymous):

You have to use all 4 numbers

OpenStudy (anonymous):

OK

OpenStudy (alexandervonhumboldt2):

(8-3-3)+8+8+3+3=24

OpenStudy (anonymous):

that's not right @AlexandervonHumboldt2

OpenStudy (anonymous):

8/(3-(8/3))

OpenStudy (nurali):

8/(3-(8/3)) = 8/(1/3) = 24

OpenStudy (anonymous):

8/(1/3) = 8 * 3 = 24

OpenStudy (anonymous):

1)8/(3-(8/3)) =8/(1/3) =24

OpenStudy (anonymous):

Is that what we said?

OpenStudy (anonymous):

Yeah @Nurali Said the answer

OpenStudy (anonymous):

I thought I said the answer

OpenStudy (anonymous):

You said the first part of the answer

OpenStudy (anonymous):

Oh you wanted the answer in parts

OpenStudy (anonymous):

No i wanted the whole answer

OpenStudy (anonymous):

\[8/(3-8/3)\]

OpenStudy (anonymous):

If he helped you out, don't forget to close the question